### v2

# Over-the-Air DPD and Reciprocity Calibration in Massive MIMO and Beyond

Abstract
Non-linear transceivers and non-reciprocity of downlink and uplink channels are two major challenges in the deployment of massive multiple-input-multiple-output (MIMO) systems. We consider an over-the-air (OTA) approach for digital pre-distortion (DPD) and reciprocity calibration to jointly address these issues. In particular, we consider a memory-less non-linearity model for the base station (BS) transmitters, and we propose a method to linearize the transmitters and perform the calibration by using mutual coupling OTA measurements between BS antennas. We show that, by using only the OTA-based data, we can linearize the transmitters and design the calibration to compensate for both the non-linearity and non-reciprocity of BS transceivers effectively. This allows to alleviate the requirement to have dedicated hardware modules for transceiver characterization. Moreover, the proposed reciprocity calibration method is solely based on closed-form linear transformations, achieving a significant complexity reduction over state-of-the-art methods, which usually rely on costly iterative computations. Simulation results showcase the potential of our approach in terms of the calibration matrix estimation error and downlink data-rates when applying zero-forcing (ZF) precoding after using our OTA-based DPD and reciprocity calibration method.
Index Terms:
Digital Pre-Distortion, Massive MIMO, Over-the-Air, Reciprocity Calibration.
I Introduction
Massive multiple-input multiple-output (MIMO) has been one of the main technologies in the development of the fifth-generation (5G) of wireless networks, by enabling significant improvements in network capacity and reliability [1, 2]. In the early stages of massive MIMO development, several proposals motivated the adoption of frequency-division duplexing (FDD) in massive MIMO deployments. While some advantages may arise from considering FDD [3], the over-head in downlink channel estimation is an important drawback which limits the system scalability. Therefore, time-division duplexing (TDD) is selected as the more viable approach for the deployment of massive MIMO in 5G and beyond, since it enables downlink channel estimation based on uplink channel state information (CSI) and channel reciprocity [4].
[…]
In this paper, we propose a method exploiting OTA measurements of inter-antenna mutual couplings at the BS to perform both the DPD and reciprocity calibration. The reciprocity methods in the literature mostly rely on linear transceivers, i.e., modeled by a complex number [6, 5]. This assumption is not accurate in practical cases, especially when scaling up massive MIMO systems, which necessitates deploying less expensive non-linear components and non-ideal linearization techniques for cost-efficiency reasons. Therefore, we assume that the TX-chains in the BS are non-linear and we propose a method to linearize them based on mutual coupling measurements. Our approach relies on the same hardware available for OTA-based reciprocity calibration, which improves resource efficiency by eliminating the need for dedicated hardware to perform per-antenna DPD. Considering the linearized transmitters after applying our OTA-based method, we derive a reciprocity calibration approach which depends only on linear closed-form transformations. Thus, the proposed reciprocity calibration allows for a significant complexity reduction over iterative methods as the ones considered in [6, 5]. Numerical results show that, although we consider non-linear TX-chains, the calibration matrix estimation error is fairly close to the Cramer-Rao Lower Bound (CRLB) derived in [6]. We also evaluate the system performance in downlink data transmission using zero-forzing (ZF) precoding and show that the proposed calibration method can approach the optimal calibration performance for moderate values of signal to noise ratio (SNR).
II System Model
We consider a TDD multi-user-MIMO scenario where an $M$-antenna BS serves $K\leq M$ user equipments through a narrow-band channel. We assume that the UE transceivers operate in the linear regime during UL and DL. The BS receivers are also assumed to operate in the linear regime for the UL, but the BS transmitters exhibit non-linear response for the DL. [^1] The assumption of having non-linear behavior only in the BS transmitters is reasonable taking into account that this is where the input power is significantly higher, pushing the power amplifiers to the non-linear regime [11].
II-A Uplink
The $M\times 1$ vector of received symbols at the BS during an UL transmission may be expressed as
$\boldsymbol{y}_{\mathrm{B}}=\boldsymbol{H}_{\mathrm{UL}}\boldsymbol{s}_{\mathrm{U}}+\boldsymbol{n}_{\mathrm{B}},$ (1)
where $\boldsymbol{s}_{\mathrm{U}}$ is the $K\times 1$ vector of input symbols to each UE TX-chain and $\boldsymbol{n}_{\mathrm{B}}\sim\mathcal{CN}(\boldsymbol{0},N_{0,\mathrm{B}}\boldsymbol{I}_{M})$ models the additive white Gaussian noise (AWGN) at the BS. The $M\times K$ channel matrix, is given by
$\boldsymbol{H}_{\mathrm{UL}}=\boldsymbol{R}_{\mathrm{B}}\boldsymbol{H}\boldsymbol{T}_{\mathrm{U}},$ (2)
where $\boldsymbol{R}_{\mathrm{B}}=\mathrm{diag}(r^{\mathrm{B}}_{1},\dots,r^{\mathrm{B}}_{M})$ and $\boldsymbol{T}_{\mathrm{U}}=\mathrm{diag}(t_{1}^{\mathrm{U}},\dots,t^{\mathrm{U}}_{K})$ are associated with the linear response of the BS receivers and the UE transmitters, respectively, and $\boldsymbol{H}$ corresponds to the $M\times K$ reciprocal propagation channel matrix [6]. Note that, if the UEs transmitters had non-linear behavior, the term $\boldsymbol{T}_{\mathrm{U}}$ in the estimated UL channel would be substituted by a non-linear function of the pilot matrix. A thorough study of its implications may be studied in future work, but the presented method would still be able to cope with the non-linearity and non-reciprocity originated at the BS side.
The UL channel, which includes the effects of the UE transmitters and the BS receivers, is assumed to be perfectly estimated at the BS based on UL pilots transmitted by the UEs. Hence, perfect knowledge of $\boldsymbol{H}_{\mathrm{UL}}$ is available at the BS, allowing for effective implementation of linear processing techniques, e.g., ZF, linear minimum mean squared error (LMMSE), and maximum ratio combining (MRC), to estimate $\boldsymbol{s}_{\mathrm{U}}$. However, the diagonal entries of $\boldsymbol{T}_{\mathrm{U}}$ and $\boldsymbol{R}_{\mathrm{B}}$, which may take arbitrary complex numbers, are unknown since they can be affected by parameters such as temperature, hardware imperfections, etc.
II-B Downlink
During the DL transmission phase, the $K\times 1$ vector of symbols received at the UEs may be expressed as
$\boldsymbol{y}_{\mathrm{U}}=\boldsymbol{R}_{\mathrm{U}}\boldsymbol{H}^{\mathrm{T}}\boldsymbol{f}(\boldsymbol{x}_{\mathrm{B}})+\boldsymbol{n}_{\mathrm{U}},$ (3)
where $\boldsymbol{x}_{\mathrm{B}}$ is the $M\times 1$ vector of input symbols to each BS TX-chain, $\boldsymbol{n}_{\mathrm{U}}\sim\mathcal{CN}(\boldsymbol{0},\mathrm{diag}(N_{0,\mathrm{U}_{1}},\dots,N_{0,\mathrm{U}_{K}}))$ models the AWGN at the UEs, $\boldsymbol{R}_{\mathrm{U}}=\mathrm{diag}((r^{\mathrm{U}}_{1},\dots,r^{\mathrm{U}}_{K})$ is associated with the linear response of the UE receivers, while $\boldsymbol{f}:\mathbb{C}^{M\times 1}\rightarrow\mathbb{C}^{M\times 1}$ is a vector-valued function modeling the non-linear response of the BS TX-chains. We assume that the transmitted symbols are generated such that
$\boldsymbol{x}_{\mathrm{B}}=\boldsymbol{g}(\boldsymbol{W}\boldsymbol{s}_{\mathrm{B}}),$ (4)
where $\boldsymbol{s}_{\mathrm{B}}$ is the $K\times 1$ vector of symbols intended for the UEs, $\boldsymbol{W}$ is the $M\times K$ linear precoding matrix applied at the BS baseband unit (BBU), while $\boldsymbol{g}:\mathbb{C}^{M\times 1}\rightarrow\mathbb{C}^{M\times 1}$ is the non-linear vector-valued function associated to the DPD applied at each TX-chain.
Let us assume that the cross-talk between TX-chains is negligible so that $\boldsymbol{f}(\cdot)$, and correspondingly $\boldsymbol{g}(\cdot)$, are component-wise functions. Considering a third-order memory-less polynomial model [12], we express
$f_{m}(\boldsymbol{x})=t^{\mathrm{B}}_{m}x_{m}+\beta_{m}x_{m}|x_{m}|^{2},\;\;\;\;\forall m\in\{1,\dots,M\},$ (5)
where $t^{\mathrm{B}}_{m}$ and $\beta_{m}$ are two complex scalars characterizing the non-linear response of the $m$’th TX-chain at the BS. The reason for focusing on the third-order non-linearity model is that it allows capturing the main idea of our solution, while maintaining a neat exposition. However, the main results of this work have trivial extension for higher order models, as remarked in the following section. Alternatively, one can fit any RF non-linear behavior to the 3rd order model, which should still capture its main impact [13]. In general, the BS TX-chain and UEs RX-chain responses are unknown, which means that the non-linear parameters $t^{\mathrm{B}}_{m}$ and $\beta_{m}$, as well as the diagonal entries of $\boldsymbol{R}_{\mathrm{U}}$, are unknown at the BS. Note that, unlike traditional work on reciprocity calibration [6], where $\boldsymbol{f}(\cdot)$ is associated to a linear transformation $\boldsymbol{T}_{B}=\mathrm{diag}(t^{\mathrm{B}}_{1},\dots,t^{\mathrm{B}}_{M})$, we cannot hereby define an aggregated DL channel matrix due to the non-linear nature of the TX-chains.
II-C Background: OTA Reciprocity Calibration
As mentioned earlier, previous work has addressed the problem of reciprocity calibration in massive MIMO assuming BS TX-chains that operate in the linear regime [6, 5]. Under such assumptions, we may define the DL channel matrix as
$\boldsymbol{H}_{\mathrm{DL}}=\boldsymbol{R}_{\mathrm{U}}\boldsymbol{H}^{\mathrm{T}}\boldsymbol{T}_{\mathrm{B}},$ (6)
which may be also derived from the presented system model, assuming $\beta_{m}=0$ in (5). [^2] Equivalently, the reciprocity calibration problem from [6] may be obtained by assuming perfect DPD up to unknown scalars, i.e., $\boldsymbol{f}(\boldsymbol{g}(\boldsymbol{x}))=\boldsymbol{T}_{\mathrm{B}}\boldsymbol{x}$. The main goal of reciprocity calibration methods, as the one in [6], is to estimate the reciprocity matrix, defined as
$\boldsymbol{C}=\boldsymbol{T}_{\mathrm{B}}\boldsymbol{R}_{\mathrm{B}}^{-1}.$ (7)
The reason is that, if we have knowledge of $\boldsymbol{C}$, we can transform the known UL channel matrix into
$\widetilde{\boldsymbol{H}}_{\mathrm{DL}}$ $=(\boldsymbol{C}\boldsymbol{H}_{\mathrm{UL}})^{\mathrm{T}}$ (8)
$=\boldsymbol{T}_{\mathrm{U}}\boldsymbol{H}^{\mathrm{T}}\boldsymbol{T}_{\mathrm{B}}.$
Note that $\widetilde{\boldsymbol{H}}_{\mathrm{DL}}$ corresponds to $\boldsymbol{H}_{\mathrm{DL}}$ up to an unknown $K\times K$ diagonal matrix, namely $\boldsymbol{D}=\boldsymbol{T}_{\mathrm{U}}\boldsymbol{R}_{\mathrm{U}}^{-1}$, multiplied from the left. Hence, $\widetilde{\boldsymbol{H}}_{\mathrm{DL}}$ can be effectively used for linear precoding, with the only caveat that the symbols received by the UEs would end up multiplied by an unknown scalar, which has negligible impact on system performance [14]. [^3] In the practical testbed considered in [6], this issue is addressed by sending an extra DL pilot. We may thus ignore the non-reciprocity associated to the UEs hardware, modeled by $\boldsymbol{R}_{\mathrm{U}}$ and $\boldsymbol{T}_{\mathrm{U}}$, and focus on characterizing the non-reciprocity associated to the BS.
An important advantage of the OTA-based calibration methods which are based on mutual coupling measurements is that they avoid the need for dedicated hardware to characterize the linear response of each transceiver chain, and can improve the cost-efficiency of massive MIMO systems [6, 5]. Similarly, we can argue that having dedicated hardware to perform DPD may compromise the cost-efficiency of MIMO systems with increasing number of antennas, e.g., massive MIMO and beyond. Thus, we next propose a method to jointly characterize the non-linear response of the BS TX-chains, as well as the resulting reciprocity matrix, to suitably design $\boldsymbol{g}(\cdot)$ and $\boldsymbol{W}$ for effectively serving the UEs in the DL.
III OTA DPD and reciprocity calibration
Our proposed OTA-based solution may be divided into three stages:
  • First, the non-linear response of the BS TX-chains, associated to $\boldsymbol{f}(\cdot)$, is estimated based on OTA mutual coupling measurements.
  • Second, the DPD, associated to $\boldsymbol{g}(\cdot)$, is designed based on the estimated non-linear response.
  • Third, reciprocity calibration is performed based on the DPD-linearized BS TX-chains, after which effective DL precoding, associated to $\boldsymbol{W}$, would become available at the BS.
III-A OTA non-linearity characterization
In this stage each BS antenna transmits $N_{\text{dpd}}\geq 2$ inter-antenna pilot signals to estimate the non-linearity parameters. The signal received at the $j$’th antenna when the $\ell$’th pilot, $\ell\in\{1,\dots,N_{\text{dpd}}\}$, is transmitted by the $i$’th antenna may be expressed as
$y_{ij,\ell}=h_{ij}r_{j}\left(t_{i}x_{i,\ell}+\beta_{i}x_{i,\ell}|x_{i,\ell}|^{2}\right)+n_{ij,\ell},$ (9)
where $h_{ij}$ is the mutual coupling gain between antennas $i$ and $j$, which is assumed fixed and known at the BS, [^4] The coupling gains may be characterized with a single complete measurement of the antenna system using a network analyzer [6]. Thus, knowledge of these may be assumed in any MIMO-related scenario with co-located TX antennas. Nevertheless, our methods can be easily extended to the case where these coefficients are unknown, as outlined throughout this work. $x_{i,\ell}$ is the $\ell$’th pilot symbol transmitted by the $i$’th antenna, and $n_{ij,\ell}\sim\mathcal{CN}(0,N_{0})$ models the measurement noise. Note that we have removed the superscript B from the parameters $r_{j}$ and $t_{i}$ for notation convenience since, as previously reasoned, we may focus on the non-reciprocity associated to the BS.
For each pair of non-linearity parameters associated to one TX-chain, there are $M-1$ relevant DPD measurements per pilot transmission, i.e., all of those originated in the same antenna, but received at different antennas. Thus, each of these measurements would share the same $t_{i}$ and $\beta_{i}$ in (9), but they would be related to a different complex gain $r_{j}$, associated to the linear response of RX-chain from the respective receiving antenna. Since the complex gains $r_{j}$ are unknown, it is not possible to directly estimate the non-linearity parameters $t_{i}$ and $\beta_{i}$ from this dataset. However, we may combine the $M-1$ measurements by averaging them after compensating for the known mutual coupling gains, so as to reduce the uncertainty, as well as the resulting noise. The combined measurements are then given by
$\tilde{y}_{i,\ell}$ $=\frac{1}{M-1}\sum_{j\neq i}\frac{y_{ij,\ell}}{h_{ij}}$ (10)
$={q_{i}}(t_{i}x_{i,\ell}+\beta_{i}x_{i,\ell}|x_{i,\ell}|^{2})+\tilde{n}_{i,\ell},$
where the uncertainty is now captured in the unknown parameter $q_{i}$, given by
$q_{i}=\frac{1}{M-1}\sum_{j\neq i}r_{j}.$ (11)
Note that one could also explore alternative optimized combinations to the simple average in (10). For example, a weighted average could be optimized assuming a specific model for $h_{ij}$ or a concrete probability distribution for $r_{j}$, but this is out of scope for this paper and may be considered in future work. On the other hand, explicit knowledge of $h_{ij}$ could be avoided by absorbing it into $q_{i}$, as further remarked in Sec. III-C.
The $N_{\text{dpd}}\times 1$ data vector $\tilde{\boldsymbol{y}}_{i}=[\tilde{y}_{i,1},\dots,\tilde{y}_{i,N_{\text{dpd}}}]^{\mathrm{T}}$ may then be used to estimate the non-linearity parameters of each antenna up to the unknown factor $q_{i}$. Since our initial aim is to compensate the non-linear response of the TX-chains, this is still possible if we know the non-linear response up to an unknown linear factor, which would only have a linear effect after the non-linearity compensation. In this case, the DPD would be designed as if the non-linearity parameters are $\theta_{1i}=q_{i}t_{i}$ and $\theta_{2i}=q_{i}\beta_{i}$. We may thus rewrite the combined data vector as
$\bm{\tilde{y}}_{i}=\bm{\Phi}_{i}\bm{\theta}_{i}+\tilde{\bm{n}}_{i},$ (12)
where $\bm{\theta}_{i}=[\theta_{1i},\theta_{2i}]^{\mathrm{T}}$ is the $2\times 1$ vector of parameters to be estimated, $\bm{\Phi}_{i}$ is the $N_{\text{dpd}}\times 2$ known pilot matrix whose columns are given by $\bm{\Phi}_{i,1}=[x_{i,1},\dots,x_{i,N_{\text{dpd}}}]^{\mathrm{T}}$ and $\bm{\Phi}_{i,2}=[x_{i,1}|x_{i,1}|^{2},\dots,x_{i,N_{\text{dpd}}}|x_{i,N_{\text{dpd}}}|^{2}]^{\mathrm{T}}$, and $\tilde{\bm{n}}_{i}\sim\mathcal{CN}(\boldsymbol{0},\varsigma_{i}\boldsymbol{I}_{N_{\text{dpd}}})$ is the resulting noise vector where
$\varsigma_{i}=\frac{N_{0}}{(M-1)^{2}}\sum_{j\neq i}\frac{1}{|{h_{ij}}|^{2}}.$ (13)
Since the noise vector is white i.i.d Gaussian, the least-squares (LS) estimator is also the minimum-variance unbiased (MVU) estimator [15], and can be used to estimate the scaled non-linearity parameters as
$\hat{\boldsymbol{\theta}}_{i}=(\bm{\Phi}_{i}^{H}\bm{\Phi}_{i})^{-1}\bm{\Phi}_{i}^{H}\bm{\tilde{y}}_{i}.$ (14)
Note that, in case of considering non-linearity polynomial models of higher order, the vector $\bm{\theta}$ (correspondingly $\bm{\Phi}$) would include one term per polynomial coefficient [10]. However, the presented method would still be applicable.
III-B DPD linearization
In this stage the non-linearity parameters estimated in the previous stage are used to linearize the output via DPD, i.e., by adjusting $\boldsymbol{g}(\cdot)$ in (4). Let us assume that $\bm{\theta}_{i}$ has been perfectly estimated from (12). The true non-linearity to compensate is the 3rd-order nonlinear function given in (5). However, the estimated non-linearity parameters in (14) characterize a different component-wise function given by
$\tilde{f}_{m}(\boldsymbol{x})$ $=\theta_{1m}x_{m}+\theta_{2m}x_{m}|x_{m}|^{2}$ (15)
$=q_{m}f_{m}(\boldsymbol{x}).$
We may thus express
$\boldsymbol{f}(\boldsymbol{x})=\boldsymbol{Q}^{-1}\tilde{\boldsymbol{f}}(\boldsymbol{x}),$ (16)
where $\boldsymbol{Q}=\mathrm{diag}(q_{1},\dots,q_{M})$.
Since the function $\tilde{\boldsymbol{f}}(\cdot)$ is fully characterized, we can find its inverse by using methods such as the postdistortion approach [16]. We may then select
$\boldsymbol{g}(\cdot)=\tilde{\boldsymbol{f}}^{-1}(\cdot),$ (17)
which is applied to the the transmitted symbols as described in (4). The resulting symbols transmitted through the reciprocal channel, which are given by $\boldsymbol{f}(\boldsymbol{x}_{\mathrm{B}})$ in (3), may then be expressed as
$\boldsymbol{f}(\boldsymbol{x}_{\mathrm{B}})$ $=\boldsymbol{Q}^{-1}\tilde{\boldsymbol{f}}(\boldsymbol{g}(\boldsymbol{W}\boldsymbol{s}_{\mathrm{B}}))$ (18)
$=\boldsymbol{Q}^{-1}\boldsymbol{W}\boldsymbol{s}_{\mathrm{B}}.$
Hence, applying the proposed OTA-DPD, results in an equivalent linear transmitter gain given by $\tilde{\boldsymbol{T}}_{\mathrm{B}}=\boldsymbol{Q}^{-1}$. Now that the transmitter is linear with an unknown complex gain, we can define a DL channel matrix equivalent to (6), but substituting $\boldsymbol{T}_{\mathrm{B}}$ for $\tilde{\boldsymbol{T}}_{\mathrm{B}}$, so that reciprocity calibration methods as those presented in [6, 5] are directly applicable. However, we will show that the reciprocity calibration can be performed without the need for complex iterative methods.
III-C Reciprocity calibration
The last stage consists of performing OTA reciprocity calibration considering the TX-chains previously linearized through the DPD linearization stage. To this end, each BS antenna transmits inter-antenna pilots to other antennas. The received symbols at the $j$’th antenna from the $i$’th antenna may be expressed as
${y}_{ij}=h_{ij}r_{j}\tilde{t}_{i}{x}_{ij}+n_{ij},$ (19)
where the variables have direct correspondence with those defined in (9), but substituting $t_{i}$ for $\tilde{t}_{i}=1/{q_{i}}$ and having $\beta_{i}=0$. The measurements defined in (19) can be directly employed to estimate the product of unknown parameters $r_{j}\tilde{t}_{i}$. In fact, we may now use the trivial MVU estimator, given by
$\widehat{r_{j}\tilde{t}_{i}}=\frac{1}{h_{ij}{x}_{ij}}{y}_{ij}.$ (20)
However, in order to perform reciprocity calibration, we are actually interested in the reciprocity parameters, $c_{m}={\tilde{t}_{m}}/{r_{m}}$, which define the adjusted reciprocity matrix entries from (7).
In [6], it was noted that multiplying all the reciprocity parameters by a common scalar does not compromise the effectiveness of the reciprocity calibration. [^5] Note that this constant may be absorbed in the linear response of the UEs RXs-chain, given by $\boldsymbol{R}_{\mathrm{U}}$ in (3). Thus, we may select one of the calibration parameters, e.g., $c_{1}$, and normalize all the rest by that value. The resulting scaled calibration parameters may then be expressed as
$\tilde{c}_{m}$ $\triangleq\frac{c_{m}}{c_{1}}=\frac{r_{1}\tilde{t}_{m}}{r_{m}\tilde{t}_{1}}$ (21)
which corresponds to the ratio of $r_{j}\tilde{t}_{i}$ products appearing in (20) for $(i,j)\in\{(1,m),(m,1)\}$. Since each of these products can be estimated through (20), we can find estimates for the scaled calibration parameters by
$\widehat{\tilde{c}_{m}}=\frac{\widehat{r_{1}\tilde{t}_{m}}}{\widehat{r_{m}\tilde{t}_{1}}}.$ (22)
The estimation error of $\widehat{\tilde{c}_{i}}$ can be reduced by averaging several estimates of $r_{1}\tilde{t}_{i}$ and $r_{i}\tilde{t}_{1}$, which is possible if each BS antenna transmits $N_{\text{cal}}\geq 2$ pilots in (19). Note further that, assuming reciprocity of the mutual coupling coefficients, i.e. $h_{ij}=h_{ji}$, we could still estimate $\tilde{c}_{m}$ using (19) without explicit knowledge of $h_{ij}$ since the coefficients would cancel each other in (22).
We have thus shown that we can estimate all the entries of the calibration matrix up to a constant, i.e., we can estimate $\tilde{\boldsymbol{C}}=\frac{1}{c_{1}}\boldsymbol{C}$, by means of simple linear estimators. This allows us to achieve reciprocity calibration without the need for high complexity iterative algorithms, such as the algorithms used in [6].
IV Numerical results
In this section, we perform simulations to validate the feasibility and assess the performance of the proposed OTA-based DPD and calibration method. The number of BS antennas and the number of single-antenna UEs are $M=100$ and $K=10$, respectively. For the BS TX-chains non-linearity parameters in (5), we fit a 3rd order polynomial to the measurement data from [13] for a Gallium Nitride (GaN) amplifier operating at 2.1 GHz at a sample rate of 200 MHz and a signal bandwidth of 40 MHz. For the RX-chains complex gains, we use the values in [6] given by $r_{m}^{B}=0.9+0.2\frac{M-m}{M}\exp(j2\pi m/M)$. To perform the DPD, we generate a look-up table based on the OTA data to implement the inverse function of the non-linear response. For the mutual coupling channel gains in (9), we have used the linear LS fit based on the measurements in [6]. We perform the simulations for different levels of a reference OTA link SNR, given by the RX-SNR for the link between the antennas with least mutual coupling gain.
Fig. 1 illustrates the average mean square error (MSE) of the calibration matrix estimation with our proposed OTA-based method for different levels of reference SNR. The calibration matrix is estimated after performing the OTA non-linearity characterization, which is performed with $N_{\text{dpd}}=500$ or $N_{\text{dpd}}=2000$ OTA transmissions. In the calibration step, we have considered $N_{\text{cal}}=200$ or $N_{\text{cal}}=500$ OTA transmissions. Note that transceiver characteristics are slowly-varying parameters [6], which means that large values of $N_{\text{dpd}}$ and $N_{\text{cal}}$ would still have a rather small impact on the total overhead. For comparison, we have also included the case with unlimited $N_{\text{dpd}}$, and the CRLB derived in [6], which assumes linear TX-chains. We can see that the calibration matrix estimation MSE for the proposed method is fairly close to the CRLB for [6], even though we are considering non-linear TX-chains. Note that, in order to approach the CRLB, which gives a theoretical bound on the achievable performance, [6] requires (at high SNR) an iterative algorithm of considerable complexity. Specifically, the reported complexity order in [6], which further outperforms previous state-of-the-art [5], is given by $\mathcal{O}(M^{2}N_{\text{ite}})$, where $N_{\text{ite}}$ is the number of iterations of the algorithm. On the other hand, in our reciprocity calibration method the complexity is given by $\mathcal{O}(MN_{\text{cal}})$, since it requires averaging $N_{\text{cal}}$ numbers where each of them is obtained by performing 2 multiplications for each of the $M-1$ antennas (excluding the reference one). Thus, for massive MIMO and beyond, where $M\gg 1$, our method may attain a significant complexity reduction. Moreover, the effect of reducing $N_{\text{cal}}$ only has a minor effect on the calibration matrix MSE, as seen from Fig. 1. We can also see that the MSE with the proposed method can approach the perfect OTA-DPD case for given $N_{\text{cal}}$, which corresponds to infinite $N_{\text{dpd}}$ in (14).
Figure 1: Average MSE of calibration matrix estimation using the proposed OTA-based DPD and Calibration.
Fig. 2 illustrates the CDF of DL data rate under ZF precoding, where we have generated $10^{4}$ DL channel realizations with i.i.d. Rayleigh fading channel, and the DL signal power for each UE is selected to achieve an average SNR of $10$ dB at their receivers. We have selected $N_{\text{cal}}=500$ for the OTA reciprocity calibration. For comparison, we have included two extreme cases with perfect DPD as benchmarks, one ideal case with perfect DL CSI available, and one baseline case with no calibration. To have a fair comparison with the benchmarks in terms of DPD performance, and since the results in Fig. 1 show very close MSE performance between limited-size and infinite-size OTA-DPD cases, we consider the infinite-size OTA-DPD scenario for the OTA plots. We can see that, for moderate values of the reference SNR, our proposed method can significantly improve the DL data rate, and it can approach the ideal case with perfect DL CSI without requiring high-complexity iterative algorithms and dedicated hardware for performing DPD and reciprocity calibration.
V Conclusion
In this paper, we have proposed an OTA-based method for DPD and reciprocity calibration in massive MIMO systems and beyond. In particular, we considered a memory-less non-linearity model for the BS transmitters and proposed to linearize the transmitters and perform the calibration by using OTA measurements of the mutual coupling among the BS antennas. We showed that, by only using the OTA data, we can effectively linearize the transmitters and perform reciprocity calibration with reduced complexity over state-of-the-art. Simulation results showed promising performance of the proposed methodology, both in terms of the calibration matrix estimation error and DL data-rates when applying ZF precoding after using our OTA-based DPD and calibration method.
Figure 2: CDF of UEs DL Data Rate.